# IntGrad MT: Eliciting LLMs’ Machine Translation Capabilities with Sentence Interpolation and Gradual MT

### v2

@@ 5 Results
GPT 3.5 0 shot 97.63 91.30 71.89 89.48 81.23 68.73 44.53
3 shot 97.83↑0.20 91.65↑0.35 72.86↑0.97 90.26↑0.78 81.43↑0.20 69.28↑0.55 45.97↑1.43
IntGradDelta 98.04↑0.40 92.42↑1.12 77.54↑5.65 92.54↑3.06 84.03↑2.80 75.60↑6.86 53.57↑9.04
IntGradThr.+Delta 97.95↑0.31 91.54↑3.03 77.27↑5.38 92.08↑2.60 83.73↑2.50 74.78↑6.05 52.31↑7.78
Llama 3.1 70B 0 shot 97.33 - 79.41 - - - -
3 shot 97.48↑0.16 - 81.19↑1.78 - - - -
IntGradDelta 97.98↑0.66 - 84.45↑5.04 - - - -
IntGradThr.+Delta 97.33↑0.00 - 84.35↑4.94 - - - -
Mistral Nemo 0 shot 96.70 88.01 66.78 81.43 38.97 71.89 43.99
3 shot 97.57↑0.86 90.81↑2.79 69.71↑2.93 88.59↑7.16 40.93↑1.95 74.25↑2.36 52.60↑8.61
IntGradDelta 97.88↑1.18 91.04↑3.03 74.19↑7.41 89.38↑7.94 46.08↑7.10 78.86↑6.97 57.02↑13.03
IntGradThr.+Delta 97.84↑1.13 91.04↑3.03 73.47↑6.69 89.31↑7.88 44.53↑5.56 78.67↑6.78 56.79↑12.81
Llama 3.1 8B 0 shot 94.99 - 69.93 - - - -
3 shot 96.27↑1.27 - 73.07↑3.14 - - - -
IntGradDelta 96.42↑1.42 - 78.18↑8.26 - - - -
IntGradThr.+Delta 95.37↑0.38 - 77.33↑7.41 - - - -
For a more balanced examination of how much gains IntGrad MT brings relative to the original performance, we compared the error reduction rate across three different metrics, xCOMET, CometKiwi, and MetricX, in Figure 4. Since the DA score reflects how ‘good’ the translation is, the error for the DA score was calculated by subtracting the score from 100 ($100-DA$). The MQM score from MetricX was used directly, as it inherently represents the error. IntGrad MT consistently reduces errors in all metrics for all languages with only one exception. Compared to Table 1 the improvements to the high resource languages are more clearly visible.
Figure 4: Error reduction rate for each target language and model, calculated as the ratio of reduced error relative to the error of zero-shot translation results. Error scores from xCOMET and CometKiwi were obtained by subtracting their values from 100, while MetricX’s score was used directly.
6 Analysis
6.1 Ablation Study
@@
To find the optimal combination of strategies for the IntGrad MT, we conducted an ablation study on En-Ko translation task using ChatGPT in terms of four different dimensions: Start sentence selection strategy, Number of start sentences, MT result aggregation strategy, and Output selection strategy. In this section, we analyze the effects of each strategy by averaging the QE scores. Full results with all different combination of ablation settings can be found in Appendix G.
6.1.1 Start Sentence Selection Strategy
@@
Based on the intuition that start sentences that are similar to the source sentence will be helpful, we utilized three different metrics — SBERT similarity(Reimers & Gurevych, 2019), Levenshtein distance(Levenshtein & others, 1966), and Tree edit distance(Zhang & Shasha, 1989) — to calculate similarity between sentences, and combined them in theee different ways. The first way is Sort, which sorts sentences by multiple metrics with varying priorities. The second way, Filter, initially selects the top 10 sentences based on SBERT similarity, then sorts the selection using the other metrics. The third way, Tops, picks the top sentence based on the highest similarity scores from each metric. As shown in Table 2, the selection strategies that produced the highest scores differed across aggregation methods. We chose sorting by SBERT similarity and then tree edit distance (Sort(S-T)), as it yielded the highest average scores.
@@
Table 2: Averaged scores with different start selection strategies, start sentence numbers and aggregation strategies applied to EN-KO translation task. ‘Sort’, ‘Filter’, and ‘Tops’ denote the start sentence selection methods. Each letter in parentheses represents a similarity metric, with the order indicating the priority of these metrics. ‘L’ stands for Levenshtein distance, ‘T’ stands for tree edit distance, and ‘S’ stands for SBERT similarity. The highest average values for each axis are highlighted in bold. The highest values for each start sentence number and aggregation strategy are highlighted in underline.
@@
6.1.2 Number of Start Sentences & MT Aggregation Strategy
When selecting the start sentences, we need to decide whether to use more than one start sentence. If we choose more than one, we must aggregate each translation result generated by Gradual MT. We investigated two distinct strategies for aggregating the results of Gradual MT. The first method, referred to as Polling, selects the MT result with the highest number of duplicates, drawing inspiration from prior research on self-consistency (Wang et al., 2023). If no duplicates are found, a result is selected randomly. The second method, Prompting, involves feeding all Gradual MT results into the LLM as few-shot examples to generate the final MT output. As shown in Table 2, the prompting strategy outperforms polling by 0.49 points. Polling is even worse than using a single start sentence.
6.1.3 Output Selection Strategy.
We tested three strategies for output selection. The first strategy, ‘Thr.’, aims to minimize costs by evaluating zero-shot translation results with a QE model and applying IntGrad MT only when the QE score falls below a certain threshold. The second strategy, ‘Delta’, prioritizes maximizing performance by applying IntGrad MT first and using its output only if the QE score exceeds that of the zero-shot translation. The third strategy, ‘Thr. + Delta’, combines the first two: applying IntGrad MT when the zero-shot translation’s QE score is below a threshold, and only if IntGrad MT’s score is higher. We employed CometKiwi, a reference-free QE model, to implement these strategies. To compare them, we analyzed xCOMET scores and score changes of selected outputs, applying the optimal strategies for start selection (‘Sort(S-T)’) and aggregation (‘Prompting’). As shown in Table 3, all strategies improved overall performance, with ‘Delta’ achieving a notable gain of over 1 point in QE scores compared to zero-shot MT. The results also indicate that ‘Thr. + Delta’ reduces interpolation by more than half while maintaining comparable performance to ‘Delta’, offering an effective compromise between computational efficiency and translation quality. Results with every combination of ablation strategies are shown in Table 7 in Appendix G.
Table 3: Average xCOMET scores and score changes of selected outputs when applying the optimal strategies for start selection and aggregation in En-Ko translation. ‘Score change’ is calculated only for the adopted outputs. ‘All’ selects every output. ‘Thr.’ applies zero-shot translation first and uses IntGrad MT only when the CometKiwi score is below a threshold. ‘Delta’ selects outputs only if they outperform zero-shot translations. ‘Thr. + Delta’ combines these two strategies. We also report the number of end sentences for which interpolation and Gradual MT is executed (‘No. of Interpolated End Sents’) and the number of end sentences for which the IntGrad MT output is selected over zero-shot translation (‘No. of Selected Outputs’). The results show that ‘Thr. + Delta’ reduces the number of interpolation by more than half while maintaining nearly the same translation performance.
Output Adoption Strategy Avg. Score Avg. Score Change of Selected Outputs No. of Interpolated End Sents(%) No. of Selected Outputs (%)
@@ 6.1.2 Number of Start Sentences & MT Aggregation Strategy
Thr. 91.50 4.49 410 (45%) 410 (45%)
Delta 92.54 5.61 911 (100%) 497 (54%)
Thr. + Delta 92.08 8.50 410 (45%) 279 (31%)
Based on the results of the ablation study, we concluded that the optimal combination of strategies is as follows: for Start Sentence Selection, sorting with SBERT similarity followed by tree edit distance proved most effective(Sort(S-T); for the Number of Start Sentences, using 3 was optimal; for Aggregation, the best method was prompting; and for Output Selection, both ‘Delta’ and ‘Thr.+Delta’ were optimal. The chosen setting for the start sentence selection strategy differs from the configuration that yields the highest xCOMET score among all ablations (92.69 in Table 6), which sorts sentences by tree edit distance, followed by Levenshtein distance, and then SBERT similarity(Sort(T-L-S)). However, after considering the averaged effects across multiple tests, we opted to prioritize the method that selects start sentences by initially sorting with SBERT similarity.
6.2 Sentence Interpolation Analysis
We conducted an analysis of the interpolated sentences used in the ablation study to verify whether they effectively interpolate between the start and end sentences. The total number of sentences in interpolated paths is 12,729.
Sentence Interpolation Error Rate.
We defined that a situation where the first and last sentences in the interpolation path are not exactly the same as the start and end sentences as an error, and executed zero-shot translation for that end sentence. Sentence interpolation was successfully executed without error in 93.68% of cases.
Progresses of Interpolation.
We examined whether LLMs genuinely interpolate between start and end sentences or simply generate random sentences. To assess this, we use SBERT to embed the interpolated sentences and calculate their Euclidean distances from the end sentence. If these distances generally decrease, it indicates successful interpolation. To mesure this, we defined progress as follows:
$\text{progress}=d_{n-1,e}-d_{n,e}$
where $d_{n,e}$ denotes the euclidian distance between nth interpolated sentence in each interpolation path and end sentence.
As shown in Figure 5, the progress of interpolated sentences is generally positive, indicating that the interpolation effectively bridges the two sentences. The average progress for all interpolated sentences is 0.14, with a standard deviation of 0.23.
We also conducted a qualitative analysis to identify patterns in interpolation. First, we sampled 100 interpolation paths for En-Ko translation containing more than three sentences and obtained sentence embeddings using SBERT. Next, we plotted each interpolation path on a 2D plane using Principal Component Analysis (PCA). Despite the variety of patterns, we identified four notable ones, which we named based on their shapes: Arc, Triangle, Zig-Zag, and Leap. Figure 5 illustrates these four major patterns. The Arc pattern was the most common. In this pattern, the sentences gradually shift toward the end sentence, following an arc-shaped trajectory. While the Arc pattern presents a relatively smooth shape, the Zig-Zag and Triangle patterns exhibit more spiky trajectories, though the sentences still progress toward the end sentence. The final pattern, Leap, shows a single large shift toward the end sentence, without the gradual progression seen in the other patterns. Although the Leap pattern might initially be perceived as a “bad interpolation,” closer examination reveals that similar intervals also appear in other patterns. Overall, our qualitative analysis showed that, regardless of the pattern, sentence interpolation typically shifts the start sentence toward the end sentence. Additional plot samples can be found in Appendix F.
Figure 5: Distribution of progress in sentence interpolation (left) and 2D scatter plots showing four major patterns of interpolation paths (right). We analyzed interpolation paths from the En-Ko ablation study. Progress indicates how much each interpolated sentence moves closer to the target sentence, with an average progress of 0.14. Scatter plots are projected from SBERT embeddings onto a 2D plane using PCA. The X and Y axes of each plot represent the first and second principal components, respectively. Arrows in each plot show the trajectory of sentence shifts from the start sentence (blue-colored dot) to the end sentence (yellow-colored dot).

### v3

@@ 4.1 Setup
GPT 3.5 Zero-shot 97.63 91.30 71.89 89.48 81.23 68.73 44.53
MAPS 97.68 92.06 - - - - -
QA Decode 97.45 91.63 75.45 90.19 78.26 73.90 45.82
BridG Post 98.04 92.63 77.90 92.57 83.65 75.41 53.11
BridG Pre & Post 97.86 92.35 77.63 92.07 83.34 75.14 51.51
Llama 3.1 70B Zero-shot 97.33 - 79.41 - - - -
MAPS 96.53 - - - - - -
QA Decode 97.37 - - - - - -
BridG Post 97.93 - 84.45 - - - -
BridG Pre & Post 97.73 - 84.31 - - - -
Llama 3.1 8B Zero-shot 94.99 - 69.93 - - - -
MAPS 79.10 - - - - - -
QA Decode 88.08 - 77.67 - - - -
BridG Post 97.25 - 78.30 - - - -
BridG Pre & Post 96.99 - 77.37 - - - -
Mistral Nemo Zero-shot 96.70 88.01 66.78 81.43 38.97 71.89 43.99
MAPS 85.75 88.71 - - - - -
QA Decode 96.33 87.18 64.42 80.34 28.86 68.59 35.37
BridG Post 97.70 90.99 73.95 89.45 45.67 78.35 57.24
BridG Pre & Post 97.70 90.99 73.82 89.34 45.18 77.83 54.55
TowerInstruct 97.69 89.89 - 91.29 - - -
NLLB 96.21 67.88 81.00 82.20 77.17 82.70 71.83
Table 2: xCOMET scores of BridG MT across different translation models, sentence bridging models, and target languages. Scores are multiplied by 100 for readability, and the two highest scores for each MT model and language are highlighted in bold and underlined. For languages not supported by each model, the corresponding cells were left blank.
MT Model Method Bridging Model DE ZH HI KO SW BN MR
GPT 3.5 BridG MT Qwen2 72B Instruct 98.04 92.63 77.90 92.57 83.65 75.41 53.11
Llama 3.1 8B 97.97 92.38 77.24 91.85 83.62 74.24 52.27
Llama 3.2 3B 97.97 92.35 76.44 91.48 83.49 74.35 52.13
Zero-shot N/A 97.63 91.30 71.89 89.48 81.23 68.73 44.53
Llama 3.1 70B BridG MT Qwen2 72B Instruct 97.93 - 84.45 - - - -
Llama 3.1 8B 97.87 - 83.97 - - - -
Llama 3.2 3B 97.85 83.65 - - - -
Zero-shot N/A 97.33 - 79.41 - - - -
Llama 3.1 8B BridG MT Qwen2 72B Instruct 97.25 - 78.30 - - - -
Llama 3.1 8B 96.93 - 77.83 - - - -
Llama 3.2 3B 97.14 77.65 - - - -
Zero-shot N/A 94.99 - 69.93 - - - -
Mistral-Nemo BridG MT Qwen2 72b Instruct 97.70 90.99 73.95 89.45 45.67 78.35 57.24
Llama 3.1 8B 97.52 90.40 72.34 87.83 43.33 76.73 52.93
Llama 3.2 3B 97.42 90.04 71.62 86.55 42.30 75.83 51.66
Zero-shot N/A 96.70 88.01 66.78 81.43 38.97 71.89 43.99
Figure 3: xCOMET scores for few-shot translations (bars) and BridG MT translations (horizontal line). Both BridG MT and few-shot scores represent the average xCOMET scores of GPT-3.5 and Mistral-Nemo.
Baselines
Since our focus is on scenarios where providing relevant few-shot examples is not feasible, we first compare BridG MT with zero-shot baselines. We then extend the comparison to conventional few-shot translation settings, where gold translations sampled from the start sentence pool are used as demonstrations (15, 30, and 50 examples). This enables us to evaluate how well BridG MT—designed specifically for settings without accessible few-shot examples—performs relative to standard few-shot approaches. We further compare BridG MT with related methods: MAPS (He et al., 2024) [^4] https://github.com/zwhe99/MAPS-mt and Quality Aware(QA) Decode (Fernandes et al., 2022) [^5] https://github.com/deep-spin/qaware-decode. Among the target languages in our experiments, MAPS provides few-shot examples only for De and Zh in its published codebase; therefore, it was evaluated only on these languages. For QA Decode, we used MBR reranking, as it showed the best performance in their paper. Finally, to provide a broader context for comparison, we benchmark BridG MT against strong baseline models, including TowerInstruct 13B (Alves et al., 2024) and NLLB-200-3.3B [^6] https://huggingface.co/facebook/nllb-200-3.3B. Among the languages supported by TowerInstruct, only De, Zh, and Ko overlapped with our target languages; thus, experiments with TowerInstruct were conducted only on these languages.
Evaluation
We employ four different QE models for evaluation; xCOMET, MetricX(Juraska et al., 2023), CometKiwi, and BLEURT(Pu et al., 2021). xCOMET, CometKiwi and BLEURT predict DA score, and MetricX predicts MQM score. MQM score assesses translation erorrs on a scale from 0 to 25, where lower scores indicate higher quality. Lastly, we report BLEU and ChrF.
5 Results
We first conducted a series of ablation experiments for En-Ko translation using GPT 3.5 to identify the optimal configuration, including start selection (Step 1) and result aggregation (Step 3). This optimal configuration was then applied in the main experiments across different translation models, sentence bridging models, and target languages. A detailed explanation of the ablation study is provided in Section 6. In this section, we present the xCOMET results of BridG MT across various LLMs and target languages. The full results are provided in Appendix F.
5.1 Comparing with Zeroshot Baselines
As shown in Table 1, BridG MT consistently improves the translation performance of LLMs, particularly for low-resource languages. For example, the performance gaps of GPT 3.5 between BridG MT and the baseline are 0.4, 1.3, 6.0, 3.1, 2.4, 6.7, and 8.6 for De, Zh, Hi, Ko, Sw, Bn, and Mr, respectively. This trend remains consistent across other evaluation metrics (Tables 12, 13, 14, 15, and 16). BridG MT also outperforms alternative methods—MAPS and QA Decode. Notably, similar patterns are observed even when using QE models other than COMET and when varying the choice of bridging model. As shown in Table 2, using smaller bridging models leads to a slight decline in performance, but the overall improvement over the baseline remains substantial.
5.2 Comparing with Baselines Using Labelled Data
We further compared our results to a few-shot setup. To provide few-shot examples under conditions similar to our approach, we built a few-shot pool by selecting 100 sentences from the FLORES-200 dev set that the model translates well—similar to how we created the start sents pool. However, in this case, we employed xCOMET, a reference-based QE model. Figure 3 presents the average translation scores of GPT-3.5 and Mistral-Nemo. The results indicate that our method is comparable to, and often outperforms, few-shot translations. This demonstrates that even without high-quality gold translations, BridG MT can surpass translations generated by an LLM equipped with gold few-shot references. Additionally, we observe a performance drop when the model is given 50-shot examples, suggesting that BridG MT achieves performance levels that cannot be matched simply by increasing the number of few-shot examples. These findings remain consistent when using Llama models and evaluating with other metrics. See Appendix F for detailed results.
6 Ablation Study
@@
To find the optimal configuration for the BridG MT, we conducted an ablation study on En-Ko translation task using GPT-3.5 in terms of four different dimensions: Start sentence selection criteria, Number of start sentences, MT result aggregation strategy, and Filtering strategy. We utilized xCOMET with gold translations in the ablation study to construct the highest-quality start pool possible. Full results with all different combination of ablation settings can be found in Appendix G.
6.1 Start Sentence Selection Strategy
@@
Based on the intuition that start sentences similar to the source sentence will be helpful, we employed three different metrics—SBERT similarity (Reimers and Gurevych, 2019), Levenshtein distance (Levenshtein and others, 1966), and tree edit distance (Zhang and Shasha, 1989)—to calculate the similarity between sentences, and we combined them in three different ways. The first approach, Sort, sorts sentences using multiple metrics with varying priorities. The second approach, Filter, initially selects the top 10 sentences based on SBERT similarity and then sorts that selection using the other metrics. The third approach, Tops, picks the top sentence based on the highest similarity scores from each metric. As shown in Table 3, the selection strategies that produced the highest scores differed across aggregation methods. We chose to sort by SBERT similarity and then by tree edit distance (Sort(S-T)), as this approach yielded the highest average scores.
@@
Table 3: Averaged scores with different start selection strategies, start sentence numbers and aggregation strategies applied to EN-KO translation task. ‘Sort’, ‘Filter’, and ‘Tops’ denote the start sentence selection methods. Each letter in parentheses represents a similarity metric, with the order indicating the priority of these metrics. ‘L’, ‘T’ and ‘S’ stands for Levenshtein distance, Tree edit distance, and SBERT similarity, respectively. The highest average values for each axis are highlighted in bold. The highest values for each start sentence number and aggregation strategy are highlighted in underline.
@@
Table 4: Average xCOMET scores and score changes of selected outputs when applying the optimal strategies for start selection and aggregation in En-Ko translation. ‘Score change’ is calculated only for the adopted outputs. We also report the number of end sentences for which bridging and Gradual MT is executed (‘# Bridging’) and the number of end sentences for which the BridG MT output is selected over zero-shot translation (‘# Selected Outputs’). The results show that ‘Pre & Post’ reduces the number of bridging by more than half while maintaining nearly the same translation performance.
Filtering Strategy Avg. Score Avg. Score Change of Selected Outputs # Bridging # Selected Outputs
@@
Pre 91.50 4.49 410 (45%) 410 (45%)
Post 92.54 5.61 911 (100%) 497 (54%)
Pre & Post 92.08 8.50 410 (45%) 279 (31%)
6.2 Number of Start Sentences & MT Aggregation Strategy
When selecting the start sentences, we need to decide whether to use more than one start sentence. If we choose more than one, we must aggregate each translation result generated by Gradual MT. We investigated two distinct strategies for aggregating the results of Gradual MT. The first method, referred to as Polling, selects the MT result with the highest number of duplicates, drawing inspiration from prior research on self-consistency (Wang et al., 2023). If no duplicates are found, a result is selected randomly. The second method, Prompting, involves feeding all Gradual MT results into the LLM as few-shot examples to generate the final MT output. As shown in Table 3, the prompting strategy outperforms polling by 0.49 points. Polling is often worse than using a single start sentence.
6.3 Filtering Strategy
We tested three strategies for filtering. The first strategy, ‘Pre-filtering’, aims to minimize costs by evaluating zero-shot translation results with a QE model and applying BridG MT only when necessary (i.e., when the QE score falls below a certain threshold). The second strategy, ‘Post-filtering’, prioritizes maximizing performance by applying BridG MT first and using its output only if the QE score exceeds that of the zero-shot translation. The third strategy, ‘Pre- & Post-filtering’ combines the first two: applying BridG MT when the zero-shot translation’s QE score is below a threshold, and only if BridG MT’s score is higher. We employed CometKiwi, a reference-free QE model, to implement these strategies. To compare them, we analyzed xCOMET scores and score changes of selected outputs, applying the optimal strategies for start selection (Sort(S-T)) and aggregation (Prompting). As shown in Table 4, all strategies improved overall performance, with ‘Post-filtering’ achieving a notable gain of over 1 point in QE scores compared to zero-shot MT. The results also indicate that ‘Pre- & Post-filtering’ reduces bridging by more than half while maintaining comparable performance to ‘Post-filtering’, offering an effective compromise between computational efficiency and translation quality. Results with every combination of ablation strategies are shown in Table 18.
6.4 Optimal Configuration
Based on our ablation study, we found the optimal strategy to be selecting start sentences using SBERT similarity and tree edit distance (Sort(S-T)), using three start sentences, aggregating Gradual MT results by feeding them as few-shot examples for the final translation, and applying post-filtering. However, given the computational cost, using both pre- and post-filtering can also be a good compromise between efficiency and accuracy.
7 Discussion
7.1 Generalizability of Start Sentence Pool
To test the robustness of the start sentence pool, we evaluated BridG MT on the NTREX-128 dataset, which has a different distribution from FLORES-200—the dataset used to construct the start pool. We used Llama-3.1-8B for sentence bridging. The results show that even when the same start pool is used to translate a different dataset, BridG MT still achieves a significant performance improvement. Detailed results are provided in Table 5.
Table 5: xCOMET results of BridG MT on the NTREX-128 dataset for various target languages.
MT Model Setting De Ch Hi Ko Sw Be Mr
GPT-3.5 Zeroshot 95.10 83.71 67.54 81.85 78.98 62.62 41.55
BridG MT 95.74 85.48 71.95 84.90 81.49 69.04 48.40
Llama-70B Zeroshot 94.39 - 75.58 - - - -
BridG MT 95.50 - 80.03 - - - -
Llama-8B Zeroshot 91.33 - 64.94 - - - -
BridG MT 93.79 - 73.47 - - - -
Mistral Nemo Zeroshot 91.02 74.36 61.13 65.10 32.99 63.66 35.19
BridG MT 94.74 80.76 68.45 78.33 39.61 72.25 47.19
7.2 Balancing Performance and Speed in BridG MT
BridG MT is useful in that it can significantly enhance the translation performance of LLMs. However, the additional inference time introduced by sentence bridging and Gradual MT can be a drawback in scenarios where fast translation is required. In such cases, it may be more important to prioritize speed, even at the cost of a slight drop in performance. We analyzed the cost-performance trade-offs of three approaches to reduce cost: using a smaller bridging model, applying bridge sampling and pre-filtering. Assuming a setting where sentences are translated one at a time, we measured the time spent on each step of BridG MT and analyzed the inference time and performance across different configurations. A detailed breakdown of the inference time and cost analysis results can be found in Appendix C.
Smaller Model for Sentence Bridging
Using smaller models can significantly reduce the time required for sentence bridging. As shown in Table 8, even with these smaller models, our approach achieves performance comparable to that of Qwen2-72B-Instruct while reducing the computation time by more than half.
Bridge Sampling
Bridge sampling is a method to reduce the computational cost by applying Gradual MT only to the first(start), middle, and last(end) sentences along the interpolation path. We tested the bridge sampling using only one start sentence. Bridge sampling reduces the time for Gradual MT by 78% while showing higher performance than baselines. Results are shown in Table 9.
Pre-Filtering
Pre-filtering reduces the overall computational cost by applying BridG MT only to sentences that the model struggles to translate well. While pre-filtering introduces an additional cost for computing QE scores, this is outweighed by the overall reduction in computational cost. According to our statistics, pre-filtering reduced the overall computation time by 28%. Results are shown in Table 10.
Results from Cost-Efficient Configuration
We further tested the cost-efficient configuration, in which we use Llama-3.1-8B for both sentence bridging and translation, and applied bridge sampling. We also applied pre-filtering, as we had already confirmed its effectiveness in our ablation study (Section 6.3). The test result on Hindi shows that applying these strategies can reduce inference time by 63%, while compromising performance by only 2.4 points.
Table 6: Comparison between the results of the main experiment (‘Main’) and a more efficient configuration (‘Efficient’). The translation model used is Llama-3.1-8B, and the target language is Hindi.
Main Efficient Zeroshot
Components Bridging Model Qwen2 72B Llama 8B -
Pre-filtering X O -
Path Sampling X O -
Post-filtering O X -
Avg. Inference Time 46.59 17.19 1.74
xCOMET 78.30 75.88 69.93
7.3 Sentence Bridging Analysis
We conducted an analysis of the sentences within the sentence bridges. In particular, we examined whether the sentences within a bridge gradually become more similar to the end sentence on both the source and target sides as illustrated in Figure 1. The analysis revealed that the embedding distance between the bridge sentences and the end sentence progressively decreases on the source side, and a similar trend is observed on the target side, where the embedding distance between the gold translation and the translations generated at each step of Gradual MT also decreases. A detailed analysis is provided in Appendix I.
7.4 Quality of a Start Pool
We analyzed the impact of the start pool quality using GPT-3.5 and Mistral-Nemo across four languages: Korean (Ko), Hindi (Hi), Swahili (Sw), and Marathi (Mr). For each language, we selected three distinct start pools from the FLORES-200 dev set, sorted by their translation xCOMET scores: High, Mid, and Low. The High start pool consists of the 100 sentences with the highest xCOMET scores. The Low start pool includes the 100 sentences with the lowest scores, while the Mid start pool is constructed using the 100 sentences located in the middle of the sorted dev set. We conducted BridG MT with a single start sentence, without applying any filtering method. The results show that the quality of the start pool had a degrading effect on the final translation when using Mistral-Nemo, whereas GPT-3.5 demonstrated robustness. Full results can be found in Appendix H.